# CQIL: Inference Latency Optimization with Concurrent Computation of Quasi-Independent Layers

### v1

@@ Abstract
The fast-growing large scale language models are delivering unprecedented performance on almost all natural language processing tasks. However, the effectiveness of large language models are reliant on an exponentially increasing number of parameters. The overwhelming computation complexity incurs a high inference latency that negatively affects user experience. Existing methods to improve inference efficiency, such as tensor parallelism and quantization, target to reduce per-layer computing latency, yet overlook the cumulative latency due to the number of layers. Recent works on reducing the cumulative latency through layer removing, however, lead to significant performance drop. Motivated by the similarity of inputs among adjacent layers, we propose to identify quasi-independent layers, which can be concurrently computed to significantly decrease inference latency. We also introduce a bypassing technique to mitigate the effect of information loss. Empirical experiments of the proposed approach on the LLaMA models confirm that Concurrent Computation of Quasi-Independent Layers (CQIL) can reduce latency by up to 48.3% on the LLaMA-33B model, while maintaining a close level of performance.
@@ 3.1 Similarity of Layer Input
While prior research Zhang et al. (2023); Din et al. (2023) has identified output similarity across layers to facilitate inference efficiency through pruning, our approach is based on a different motivation with a more systematic perspective. Experimental results of previous works indicate that layer pruning often result in substantial performance declines. For instance, although the Sheared-LLaMA-1.3B model has a similar hidden dimension size to that of the Sheared-LLaMA-2.7B, it contains eight fewer layers and consequently results in significantly performance drop on downstream tasks. [^2] During the ARR February Review period, preprints Men et al. (2024); Gromov et al. (2024) introduced methodologies for pruning layers based on input similarity. We further discuss these pruning methods in the Appendix D. Contrary to removing layers, our approach seeks to parallelize layer computation to reduce inference latency while preserving the model performance. In addition, our method is orthogonal to pruning techniques and the compatibility is further demonstrated in Section 5.6.
@@ 5.5 Comparison with Tensor Parallelism
Both tensor parallelism and our method utilize additional GPUs to reduce latency. Therefore, we assess the latency reduction ratio between these two approaches. Given that DeepSpeed inference framework involves optimizations beyond tensor parallelism, it is challenging to make a direct comparison of latency. Therefore, we first measure the latency of both methods using a single GPU under various batch sizes, then evaluate the relative latency reduction ratio with an increasing number of GPUs. The comparison is carried out on the LLaMA-13B model, which is sufficiently large for tensor parallelism to be applied across four GPUs.
@@
As shown in Figure 4, tensor parallelism fails to accelerate inference for small batch sizes. In contrast, CQIL consistently offers latency reductions for LLaMA-13B, making it more amenable to the scenario of online inference. Moreover, tensor parallelism is complementary to our approach, allowing for the parallel computation of bottom layers, which are sensitive to input substitution. Integrating tensor parallelism with our approach may reduce the latency further. Due to time constraints, we leave the integration for future work.
@@ 7 Conclusion
This work has two primary limitations. First, CQIL requires additional GPUs to reduce the inference latency, limiting its application in the environment equipped with only a single GPU. Second, not every layer in LLMs could be computed concurrently, especially for layers at the bottom and top. These layers are computed on a single GPU, leaving remaining GPUs ignored. To efficiently utilize all GPUs, we suggest to apply tensor parallelism for these layers. Due to time constraints, the integration of tensor parallelism and our approach is leaved for future work.

### v2

@@ Abstract
The fast-growing large scale language models are delivering unprecedented performance on almost all natural language processing tasks. However, the effectiveness of large language models are reliant on an exponentially increasing number of parameters. The overwhelming computation complexity incurs a high inference latency that negatively affects user experience. Existing methods to improve inference efficiency, such as tensor parallelism and quantization, target to reduce per-layer computing latency, yet overlook the cumulative latency due to the number of layers. Recent works on reducing the cumulative latency through layer removing, however, lead to significant performance drop. Motivated by the similarity of inputs among adjacent layers, we propose to identify quasi-independent layers, which can be concurrently computed to significantly decrease inference latency. We also introduce a bypassing technique to mitigate the effect of information loss. Empirical experiments of the proposed approach on the LLaMA models confirm that Concurrent Computation of Quasi-Independent Layers (CQIL) can reduce latency by up to 48.3% on LLaMA-33B, while maintaining a close level of performance.  [^1] Code is available at https://github.com/Photooon/CQIL
@@ 3.1 Similarity of Layer Input
While prior research Zhang et al. (2023); Din et al. (2023) has identified output similarity across layers to facilitate inference efficiency through pruning, our approach is based on a different motivation with a more systematic perspective. Experimental results of previous works indicate that layer pruning often result in substantial performance declines. For instance, although the Sheared-LLaMA-1.3B model has a similar hidden dimension size to that of the Sheared-LLaMA-2.7B, it contains eight fewer layers and consequently results in significantly performance drop on downstream tasks. Contrary to removing layers, our approach seeks to parallelize layer computation to reduce inference latency while preserving the model performance. In addition, our method is orthogonal to pruning techniques and the compatibility is further demonstrated in Section 5.6.
@@ 5.5 Comparison with Tensor Parallelism
Both tensor parallelism and our method utilize additional GPUs to reduce latency. Therefore, we assess the latency reduction ratio between these two approaches. Given that DeepSpeed inference framework involves optimizations beyond tensor parallelism, it is challenging to make a direct comparison of latency. Therefore, we first measure the latency of both methods using a single GPU under various batch sizes, then evaluate the relative latency reduction ratio with an increasing number of GPUs. The comparison is carried out on the LLaMA-13B model, which is sufficiently large for tensor parallelism to be applied across four GPUs. The maximum batch size of tensor parallelism is 64 due to the limit of available GPU memory.
@@
As shown in Figure 4, tensor parallelism does not significantly accelerate the inference for small batch sizes. In contrast, CQIL consistently offers latency reductions for LLaMA-13B, making it more amenable to the scenario of online inference. Moreover, tensor parallelism is complementary to our approach, allowing for the parallel computation of bottom layers, which are sensitive to input substitution. Integrating tensor parallelism with our approach may reduce the latency further. Due to time constraints, we leave the integration of tensor parallelism with CQIL for future work.
@@ 7 Conclusion
This work has three primary limitations. First, CQIL requires additional GPUs to reduce the inference latency, limiting its application in the environment equipped with only a single GPU. Second, not every layer in LLMs could be computed concurrently, especially for layers at the bottom and top. These layers are computed on a single GPU, leaving remaining GPUs ignored. To efficiently utilize all GPUs, we suggest to apply tensor parallelism for these layers. Due to time constraints, the integration of tensor parallelism and our approach is leaved for future work. Finally, the computing FLOPS remain unchanged. Therefore, from an energy consumption perspective, CQIL does not result in electricity savings.
@@ Appendix D The Potential Performance Degradation of Pruning
Table 8: Numerical values of input substitution experiments (Layer 1-16).
Layer 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16
k=1 868.2 6.4 14.1 6.1 6.0 6.0 6.0 6.0 6.0 6.0 6.0 6.0 6.0 6.0 6.0
k=3 16.3 13.7 11.2 6.2 6.2 6.1 6.1 6.1 6.1 6.1 6.1 6.1 6.1
k=7 8.8 8.9 8.3 6.3 6.3 6.3 6.2 6.2 6.2
Table 9: Numerical values of input substitution experiments (Layer 17-32).
Layer 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32
k=1 6.0 6.0 6.0 6.0 6.0 5.9 5.9 5.9 5.9 5.9 5.9 5.9 6.0 6.0 6.0 7.5
k=3 6.1 6.1 6.0 6.0 6.0 6.0 6.0 6.0 6.0 6.0 6.0 6.0 6.1 6.1 6.4 8.7
k=7 6.4 6.3 6.2 6.2 6.2 6.2 6.2 6.2 6.1 6.2 6.2 6.2 6.7 6.5 6.9 10.6
Figure 8: MMLU Sensitivity of Input Substitution. We individually replace the input of layer $l$ with that of the layer $l-k$ and evaluate the MMLU score with zero-shot setting. A darker block indicates a lower MMLU score and diminished performance. Note that the zero-shot MMLU score of the original model is 33.4.
@@ Appendix E Downstream Tasks Performance of CQIL-Models without Fine-Tuning
Appendix F Additional Details of Input Substitution Experiments
To clarify the details of Figure 2, we give the numerical values of each substitution experiment in Table 8 and 9. Moreover, we conducted the sensitivity experiment on LLaMA-7B with MMLU under zero-shot setting. The results in Table 8 are similar to Figure 2, revealing that the majority of middle layers are relatively insensitive.
Appendix G Additional Bypassing Experiments
Table 10: Effect of bypassing on LLaMA-7B with $p=2$.
Bypassing Downstream Tasks Latency
Distance ($d$) Average Score Reduction
0 62.6 27.6%
1 63.2 27.0%
We conducted additional experiments for LLaMA-7B with $p=2$. The results in Table 10 confirm that bypassing technique improves the performance on all LLaMA-7Bs with different $p$ values.